# Selective Transformer for Hyperspectral Image Classification

### v2

@@ III-D2 Performance Contribution of Spatial-Spectral Mechanism
TABLE VII: Comparison of Model Parameters (K), Inference Time (s) and OA (%) Across Four Datasets
@@
Dataset Method SpectralFormer SSFTT GAHT MorphFormer GSCViT SFormer
@@
III-E Performance Comparison and Analysis
  • 1)
    3D-CNN[29]: Spatial-spectral classification using 3D-CNN. We follow the implementation with the original article.
  • 2)
    SSFCN[26]: Spatial-spectral classification with an end-to-end fully convolutional network. The model is trained using the Adam optimizer with a learning rate of $45e-$ 4 and a momentum of 0.9.
  • 3)
    SACNet[38]: Spatial-spectral classification method that integrates self-attention learning with context encoding. The network is trained using the Adam optimizer with a learning rate of $45e-$ 4 and a weight decay of $55e-$ 5.
  • 4)
    FcontNet[36]: Multiscale classification based on non-local contextual attention across space, bands, and scales. The base learning rate is 0.01, adjusted using the poly scheduling policy. The model is optimized using SGD with a momentum of 0.9 and a weight decay of $41e-$ 4.
  • 5)
    SpectralFormer[22]: A Transformer-based architecture that employs sequential learning, focusing on adjacent bands and patches. The model is trained using the Adam optimizer with a learning rate of $45e-$ 4, which decays by a factor of 0.9 after every tenth of the total epochs.
  • 6)
    SSFTT[43]: Spatial-spectral classification method integrates a CNN with a single-layer transformer and a Gaussian distribution-weighted tokenization module. The model is trained using the Adam optimizer with a learning rate of $31e-$ 3.
  • 7)
    GAHT[48]: Spatial-spectral classification through a hierarchical transformer with grouped pixel embedding. The model is trained using the SGD optimizer with a momentum of 0.9, a weight decay of $41e-$ 4, and a learning rate of $31e-$ 3.
  • 8)
    MorphFormer[49]: A ViT-based spatial-spectral classification method that combines attention mechanisms with morphological operations. The model is trained using the Adam optimizer with a weight decay of $35e-$ 3 and a learning rate of $45e-$ 4.
  • 9)
    GSCViT[45]: A lightweight network utilizing groupwise separable convolution and ViT in each feature extraction block. The model is trained with the AdamW optimizer, featuring a learning rate of $31e-$ 3 and a weight decay coefficient of 0.05.
Figure 12: Visualization of the objects with different receptive fields selection weights from Pavia University dataset. (a) Position of three objects P1: Asphalt, P2: Meadows, and P3:Bare soil. (b) Weights selection values of different receptive fields.
III-E1 Quantitative and Qualitative Results Analysis
The quantitative results include the accuracy per class, OA, $\kappa$, and AA as shown in Table III to Table VI, with the best results highlighted in bold. Taking the Pavia University dataset as an example, CNN-based HSI classification methods such as 3D-CNN at the patch level, the classification performance is limited by the size of the input patch, resulting in an OA of only 77.01%. For image-based methods such as SSFCN, although the entire image is used as input, the limited receptive field of the convolutional layer hampers the effective extraction of contextual information, leading to a lower OA of 81.48%. The SACNet method introduces a self-attention learning mechanism and context encoding to establish long-range dependencies, but its performance improvement is restricted by spatial resolution during the self-attention mechanism calculation, yielding an OA of 83.27%. In contrast, FcontNet achieves a relatively higher accuracy with an OA of 90.54% by combining scale, spatial, and channel attention mechanisms for multiscale classification. However, the extensive self-attention computations and multiscale feature design of this method result in high computational resource demands and low operational efficiency.
For transformer-based methods, both SpectralFormer and SSFTT are built on the ViT structure and learn by constructing sequence data from adjacent bands. However, due to redundancy between bands and the lack of spatial context modeling, these methods perform slightly worse than the CNN-based HSI classification methods aforementioned, with an OA of 76. 46% and 75.79%, respectively. GAHT builds a hierarchical transformer on the ViT foundation by integrating grouped pixel embeddings, MorphFormer incorporates morphological operations into the ViT attention mechanism, and GSCViT uses groupwise separable convolution with ViT to extract local and global features. All of these methods achieve relatively high accuracy, with OAs of 90.69%, 90.6%, and 93.40%, respectively. However, these approaches overlook the varying range of background information required to identify different types of land cover in the HSI, and the traditional self-attention mechanism includes redundant and irrelevant information during computation. Furthermore, insufficient refinement in capturing the interspectral context limits the accuracy of existing transformer-based HSI classification methods.
Figure 13: Visualization of (a) the dense attention matrix and (b) selective attention mask in Pavia University dataset.
To address the aforementioned issues, we employ KSTB for adaptive receptive field selection, followed by TSTB to intelligently choose the most relevant tokens for self-attention computation. Our method, SFormer, achieved an optimal OA of 96.59%, surpassing the second-best method by 3.2 percentage points. Notably, SFormer obtained the best results in nearly all per-class accuracy metrics, except for the ninth class, shadows, where it was slightly outperformed by MorphFormer. This may be attributed to the integration of morphological operations within the attention mechanism, which likely provides more effective semantic features for shadow recognition. Similarly, our proposed method demonstrated superior accuracy in three additional datasets, highlighting its advantages and applicability.
We also present classification maps in Fig. 7-10 to evaluate various methods. Using the WHU-HongHu dataset as a case study, we observed that methods such as 3D-CNN and SSFCN exhibit significant salt-and-pepper noise. In contrast, FcontNet and the latest Transformer-based approaches yield smoother and better results. A particularly noteworthy observation is on the classification of the cotton class, where most existing methods tend to misclassify it as cotton firewood. However, our SFormer model significantly reduces these errors, producing outputs that are much closer to the ground truth. This advantage has been consistently validated across multiple datasets, highlighting the superior performance of the SFormer model in handling complex classification tasks.
III-E2 Robustness with different training samples
To assess the stability of the proposed SFormer method, we conducted extensive experiments on four diverse datasets, each with varying numbers of training samples. Specifically, we randomly selected 25, 50, 75, and 100 samples per class to construct the training sets. Fig. 11 compares the classification performance in these different sample sizes. The results indicate that the classification accuracy of SFormer is steadily improving as the number of training samples increases, demonstrating its robustness. Furthermore, SFormer consistently outperforms existing state-of-the-art methods, achieving the highest accuracy across all training sample configurations.
III-E3 Computation Cost Analysis
@@
Given that previous comparative experiments indicate that transformer-based models typically achieve higher accuracy, we performed a comprehensive comparison of all transformer-based methods regarding parameters, inference time, and OA results, as presented in Table VII. In our performance evaluation, the proposed SFormer exhibited strong competitiveness and achieved a good balance between precision and efficiency in four benchmark datasets. Regarding model parameters, our SFormer has fewer parameters than the transformer-based state-of-the-art methods GAHT and GSCViT. Although the inference time of SFormer is slightly higher than that of the comparison methods, it outperforms SpectralFormer. These results suggest that SFormer successfully achieves an optimal trade-off between computational resource usage and the ability to capture complex features.
III-F Visualization
III-F1 Different Objects of Selection Weights
@@
To validate differences in receptive field requirements for various types of land cover in the HSI classification, we present the visualization of objects with different selection weights for receptive fields. Fig. 12 (a) presents the locations of three land covers: Asphalt, Meadows and Bare Soil within the Pavia University dataset and their corresponding positions. And Fig. 12 (b) shows the weight values $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ for different receptive fields used in the final fusion within the KSTB. The weight results show that the weight corresponding to the larger receptive field in P1 is 54. 07%, which confirms our previous analysis that asphalt requires more contextual information to perform an accurate classification. In contrast, at locations P2 and P3, both of which exhibit spatial continuity, smaller receptive field characteristics are sufficient for classification needs.
III-F2 Attention Matrix Change of Token Selection
@@
To further understand the effect of the token selection operation, we present the changes in the attention matrix before and after the operation. As shown in Fig. 13, we used the Pavia University dataset with the number of groups set to 2 and a selection rate of 40%. Specifically, we selected the dense attention matrix from the first head of the multi-head self-attention mechanism. From the figure, it is evident that after the token selection operation, only 40% of the elements in the original dense attention matrix retain their values, while the remaining 60% are set to 0, resulting in a sparse selective attention mask. After deep feature extraction by the TSTB, the relevant information for classification is retained, whereas irrelevant and redundant information is discarded.
IV Conclusions
In this article, we introduce a novel SFormer for the HSI classification. The SFormer is designed to dynamically select receptive fields to capture both spatial and spectral contextual information, while reducing the influence of redundant data by emphasizing the most relevant features. Recognizing that objects in HSI of varying scales require distinct spatial-spectral contextual information, the SFormer incorporates a KSTB, which dynamically selects appropriate receptive fields for effectively spatial-spectral feature extraction. Furthermore, considering vanilla self-attention mechanisms may be susceptible to redundant and irrelevant information, which can negatively affect classification performance. We introduce a TSTB that selects the most valuable tokens with the highest attention values, rather than applying dense attention across all query-key pairs. These selective mechanisms allow SFormer to capture relevant information within the optimal receptive field, enhancing its ability to effectively extract valuable spatial-spectral features. Extensive experiments conducted on four benchmark HSI datasets demonstrate that the proposed SFormer achieves competitive performance compared to existing state-of-the-art methods, thereby validating the effectiveness of the proposed SFormer.

### v3

@@ 4.5 Performance Comparison and Analysis
Table 7: Comparison of model parameters (K), inference time (s) and OA (%) across four datasets
@@
Dataset Method SpectralFormer SSFTT GAHT MorphFormer GSCViT DSFormer
@@
(4) Computation Cost Analysis. Given that previous comparative experiments indicate that transformer-based models typically achieve higher accuracy, we performed a comprehensive comparison of these methods regarding parameters, inference time, and OA results, as presented in Table 7. In our performance evaluation, the proposed DSFormer exhibited strong competitiveness and achieved a good balance between precision and efficiency in four benchmark datasets. Regarding model parameters, our DSFormer has fewer parameters than the state-of-the-art methods GAHT and GSCViT. For inference time, DSFormer is faster than SpectralFormer while achieving higher OAs. We believe the efficiency of DSFormer is primarily influenced by the grouping strategy in TSFTB, which generates a large number of tokens for subsequent self-attention calculations. In summary, these results suggest that DSFormer enables a good trade-off between computational resource usage and complex scene perception.
Figure 13: Visualization of the alterations in the attention matrix before and after the token selective operation using the Pavia University dataset. (a) The dense attention matrix. (b) The selective attention mask. The attention map is obtained from the first head of MHSA with a group number of 2 and a selection rate of 40%.
4.6 Visualization
@@
(1) Different Objects of Selection Weights. To validate the differences in the receptive field requirements for various types of land cover in HSIs, we present visualizations of the selection weights for different objects during receptive field fusion. Fig. 12 (a) presents the locations of three land covers: Asphalt, Meadows and Bare Soil within the Pavia University dataset and their corresponding positions. And Fig. 12 (b) shows the weight values $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ in equation 7 (normalized for visualization) for different receptive fields used in the final fusion within the KSFTB. The weight results show that the weight corresponding to the larger receptive field in P1 is 54.07%, which confirms our previous analysis (see Fig. 1) that asphalt requires more contextual information to perform an accurate classification. In contrast, at locations P2 and P3, both of which exhibit spatial continuity, smaller receptive field characteristics are sufficient for classification.
@@
(2) Attention Matrix Change of Token Selection. To further understand the token selection process in TSFTB, we present the changes in the attention matrix before and after the operation (corresponding to equation 13). As shown in Fig. 13, we used the Pavia University dataset with a group number of 2 and a selection rate of 40%. Specifically, we selected the dense attention matrix from the first head of the MHSA mechanism. From the figure, it is evident that after the token selection operation, only 40% of the elements in the original dense attention matrix retain their values, while the remaining 60% are set to 0, resulting in a sparse selective attention mask for subsequent effective fusion. Through this process, the relevant information for classification is retained, whereas irrelevant and redundant information is discarded.
5 Conclusion
In this article, we introduce a novel network called DSFormer for HSI classification. DSFormer flexibly selects and fuses features across varying receptive fields, achieving contextual modeling of crucial spatial-spectral features through adaptive attention selection and fusion. Recognizing that objects in HSI of varying scales require distinct spatial-spectral contextual ranges, DSFormer incorporates a KSFTB, which enables the learning of optimal receptive fields by dynamically selecting and integrating multiscale features extracted through multiple dilated convolutions. Furthermore, since the original MHSA may introduce redundant and irrelevant regions and targets that negatively affect HSI classification performance. We develop a TSFTB that realize the adaptive selection of the most valuable spatial-spectral tokens during the context fusion process under a spectral grouping strategy. These selection and fusion mechanisms collectively enable DSFormer to achieve effective visual perception of spatial-spectral contexts within an optimal receptive field. Extensive experiments conducted on four benchmark HSI datasets demonstrate that the proposed DSFormer achieves competitive comprehension performance on accuracy, robustness and inference efficiency compared to existing advanced methods.